# Formal Analysis of the Contract Automata Runtime Environment with Uppaal: Modelling, Verification and Testing

### v1

@@ 1. Introduction
The formal model is a network of stochastic timed automata as accepted by the Uppaal toolbox. Different variants of the formal model are proposed. The model undergoes verification against the desired properties using Uppaal, employing both exhaustive model checking, statistical model checking and model-based testing. This combination allows for thorough analysis of the model’s behavior and ensures scalability when dealing with systems that possess a large state-space.
This work is an extension of the conference paper [7]. The extension includes the addition of more figures and tables, as well as further explanations throughout the paper. Testing, which was only briefly mentioned in [7], is now thoroughly described. We cover both the generation of abstract tests from the Uppaal model and the process of concretising these tests in the real application.
@@ 3.2. Uppaal
Uppaal SMC [23] is a variant of Uppaal [19], which is a well-known toolbox for the verification of real-time systems. Uppaal models are stochastic timed automata, in which non-determinism is replaced with probabilistic choices and time delays with probability distributions (uniform for bounded time and exponential for unbounded time). These automata may communicate via broadcast channels and shared variables. In this paper, we will use both exhaustive and statistical model checking. Statistical Model Checking (SMC) [37] involves running a sufficient number of (probabilistic) simulations of a system model to obtain statistical evidence (with a predefined level of statistical confidence) of the quantitative properties to be checked. Monte Carlo estimation with Chernoff-Hoeffding bound executes $N=\lceil(ln(2)-ln(\alpha))/(2\varepsilon^{2})\rceil$ simulations $\rho_{i}$, $i\!\in\!1...N$, to provide the interval $[p^{\prime}-\varepsilon,p^{\prime}+\varepsilon]$ with confidence $1-\alpha$, where $p^{\prime}=(\#\{\rho_{i}\!\mid\!\rho_{i}\!\models\!\varphi\})/N$, i.e., $\mathtt{Pr}(|p^{\prime}-p|\leq\varepsilon)\geq 1-\alpha$ where $p$ is the unknown value of the formula $\varphi$ being estimated statistically and $\varepsilon$ and $\alpha$ are the user-defined precision and confidence, respectively. Crucially, the number of simulations used to estimate a formula is independent of the model’s size and depends only on the parameters $\alpha$ and $\varepsilon$. In practice the number of simulations required by Uppaal to reach a specific confidence level is optimized and is thus lower than the above theoretical bound. Uppaal supports template automata used to instantiate different copies (in different experiments) of the same automaton, distinguishable by their parameters. The selection of Uppaal as the chosen tool is influenced by several factors. These include its extensive adoption by the community, expertise of the author, usability, primitive support for real-time and stochastic modeling, probabilistic and non-deterministic choices, and capabilities for statistical model checking, simulation, and model-based testing.
@@
The off-line test case generator of Uppaal is called Yggdrasil, which is now integrated into the Uppaal GUI as a tab. Starting from the Uppaal models, a suite of test cases is created, aiming to cover all syntactic transitions in the model (i.e., edge coverage). The Uppaal model can be decorated with test code, which can be generated from the execution of a transition or at the entry to or exit from a location. Any language can be used for the test code. An abstract test case is generated from a trace of execution of the Uppaal model. During the execution of the trace, the test code of the model is dumped into a text file, forming the abstract test case. When the system is composed of a set of automata, it is possible to generate the test code from a single automaton or from the whole system. In this latter case, the single generated abstract test case contains the interleaved test code from all automata. The traces of execution can be generated as witnesses of some reachability property defined using the Uppaal syntax. Random traces can also be executed to improve test coverage. Parameters to be set include the depth of the trace, the type of search (i.e., breadth, depth), and options for the trace to be generated (i.e., fastest, shortest). The generated abstract test cases need to be decorated and filled with the missing information from the system under test to become concrete test cases.
@@ 3.3. CARE
The Contract Automata Runtime Environment (CARE[11] provides facilities for pairing the contract automata specifications with actual implementations of service-based applications. Note that our purpose is not to formally analyse the applications created using CARE, but to formally analyse CARE itself. The formal verification of applications developed using CARE is a consequence of the reliance of CARE on the formal guarantees provided by contract automata [11], along with the formal verification of CARE as an application. This paper addresses the latter aspect. This software is organised into classes for the orchestrated services (cf. Figure 1) and classes for the orchestrator.The two core abstract Java classes are the RunnableOrchestration and the RunnableOrchestratedContract.
@@
Two aspects to implement are choices and termination (through the abstract method choice). CARE is equipped with default implementations, but can be extended (by implementing the relative interfaces and abstract methods) to include other options, other than the default ones. Currently, a so-called ‘dictatorial’ choice (i.e., an internal choice of the orchestrator, external for the services) and a so-called ‘majoritarian’ choice (services vote and the majority wins) are two implemented options. MajoritarianChoiceRunnableOrchestratedContract and DictatorialChoiceRunnableOrchestratedContract are the two classes specialising RunnableOrchestratedContract according to how the choice is handled and implementing the abstract methods. CARE also provides default implementations for the low-level message exchanges. Currently, the two available options are the ‘centralised’ action, where the orchestrator acts as a proxy, and the ‘distributed’ action, where two services matching their actions directly interact with each other once the orchestrator has made them aware of a matching partner and its address/port. Accordingly, each RunnableOrchestratedContract has an OrchestratedAction (instance variable act) used to implement the corresponding actions that can be either distributed or centralised according to the current implementation.
@@
The class diagram of the orchestrator side in Figure 2. The abstract class RunnableOrchestration implements a special service that reads the synthesised orchestration (stored in the instance variable contract) and orchestrates the services (the instances of RunnableOrchestratedContract) to realise the overall application. The instance variables port and addresses contain the corresponding addresses and ports of the services involved in the orchestration. Similarly to the case of the orchestrated contract, also the orchestrator is specialised according to either a dictatorial or a majoritarian implementation of the abstract method choice. Moreover, an OrchestratorAction instance variable is used to implement each action of the orchestration, either centralised or distributed, thus matching the corresponding actions of the orchestrated services.
@@ 4. Methodology
In this section, we describe both the methodology used for modeling the communications, abstracting away irrelevant details, and the adequacy of the model to the actual implementation.
@@ 4.1. Modelling TCP/IP sockets communication
The two global (FIFO) arrays in the model are named orc2services[N][queueSize] and services2orc[N][queueSize] (see Figure 3 left). These are used to model the TCP/IP sockets buffers, where N is the number of services in the model and queueSize is the size of the queue of the components. These arrays are only modified by functions for enqueueing and dequeuing messages.
@@
The local declarations of the two automata contain methods for sending and receiving from the partners and checking their queues of messages. Both automata declare a method enqueue for sending to the partner a message. Messages are modeled as global integer constants, depicted in Figure 3 left. Indeed, in this model the actual payload of each communication is abstracted away The enqueue method of the services only takes as parameter the signal ide_sig to send (see Figure 3 center). The identifier of the service id is only needed on the orchestrator side to identify the partner (there is only one orchestrator thus no identifier is needed for it, see Figure 3 right). Similarly, both automata have a method dequeue for consuming messages from their respective arrays.
@@
Source locations of sending and receiving transitions are neither committed nor urgent. In fact, an enabled committed transition (i.e., whose source location is committed, denoted with C) must be executed before any other non-committed transition in the network. Instead, an urgent transition must be executed without any delay. If a send or receiving transition were to be either committed or urgent, it would introduce the possibility of false positive deadlocks. This scenario arises when, for example, the receiver has a full buffer (i.e., array) and is prepared to free it, but the sender transition is enabled and committed. Similarly, this false positive can occur when the buffer is full, the send transition is urgent but the receive operation is not urgent, or vice versa, when the buffer is empty, the read transition is urgent, and the send operation is not urgent.
Hence, the operations of writing to and reading from a buffer are represented using stochastic delays, specifically following an exponential distribution. Two rates are employed to capture the delay associated with reading and writing. These exponential delays are present in all non-committed and non-final locations of both automata in Figure 6 and Figure 7. However, the presence of unbounded delays introduces scenarios where the receiver (resp., sender) may wait indefinitely without executing its read (resp., write) transition, even when it is enabled. These scenarios would invalidate the exhaustive model checking of reachability properties that are satisfied by the actual system, leading to false positives. In the real implementation, Java TCP/IP sockets offer a timeout mechanism wherein an exception is thrown if no message exchange occurs within a specified time frame. All sockets used by CARE have this timeout. Consequently, a dedicated automaton called SocketTimeout is replicated for each service to model the timeout operation (see Figure 8 left). Each send or receive operation in every socket resets the SocketTimeout clock c. If no reset operation is received within a certain duration (variable timeout), SocketTimeout enters a location called Timeout and broadcasts the signal fail, indicating that an exception has been thrown. All automata have a transition from every location (except for Terminated) to a location Timeout that is reached upon receiving the signal fail. For readability, these Timeout locations and all their incoming transitions are not shown in Figure 6 and Figure 7.
@@ 4.2. Abstractions
We have discussed the modelling of Java TCP/IP socket communications. We now discuss other aspects that have been abstracted away in the model. We note that the abstracted aspects are irrelevant for the analysis discussed in Section 6. In the model, the underlying orchestration automaton is abstracted together with the contracts of the services. Thus, all conditionals that are dependent from the underlying orchestration are abstracted as probabilistic choices. We assume that the orchestration has been correctly synthesised from the services contracts. This allows us to verify the interactions for any possible valid orchestration. If a specific orchestration would be modelled, then we would lose such generality. In particular, the payloads of the communications (e.g., which specific action, which choices) are abstracted. The conditions used to decide whether to perform a choice, an action or to stop are also abstracted away. The only modelled condition is that no two consecutive choices are allowed (i.e., after choosing, the chosen step must be performed). When executing a transition, the conditions used to check whether the label of the transition is an offer or a match are abstracted away in the model. Moreover, the identifiers of the services involved in performing a choice or an action (which are concretely extracted from the labels of the transitions) are also chosen non-deterministically. Indeed, all services are distinguished replicas of the same automaton. Finally, we model a single orchestration. In fact, when multiple orchestrations are executed, they operate independently of each other and can be verified individually.
4.3. Adequacy
@@
Traceability and model-based testing have been used to ascertain the adequacy of the model with respect to the implementation. Traceability involves connecting the items of the abstract model with the lines of source code that are being abstractly represented. In [6] each transition of the model is traced back to the corresponding source code instructions using comments in the model. Traceability information is related to a specific version of the source code (i.e., b1a6954). For example, Figure 4 shows a portion of the automaton modelling the RunnableOrchestratedContract (discussed in details in the next section). Specifically, from state Ready a transition is executed which is traced back to line 98 of the class RunnableOrchestratedContract, also showed in Figure 4. Indeed, at line 98 the first read operation from the ObjectInputStream, reading from the TCP/IP socket, is performed by the service. This is abstracted in the model by the operation dequeue() of the automaton.

### v2

@@ 1. Introduction
The formal model is a network of stochastic timed automata as accepted by the Uppaal toolbox. Different variants of the formal model are proposed. The model undergoes verification against the desired properties using Uppaal, employing both exhaustive model checking, statistical model checking and model-based testing. This combination allows for thorough analysis of the model’s behaviour and ensures scalability when dealing with systems that possess a large state-space.
@@
This work is an extension of the conference paper [11]. The extension includes the addition of more figures and tables, as well as further explanations throughout the paper. The most significant addition concerns the testing phase, which was only briefly mentioned in [11]. Testing is now thoroughly described in Section 7. We cover both the generation of abstract tests from the Uppaal model and the process of concretising these tests in the real application.
Summarising, the core contributions of this paper are: (i) the bottom-up formal modelling of CARE, an already established, open-source distributed middleware, modelled as a network of stochastic timed automata suitable for the Uppaal toolbox; (ii) the verification of desired properties of the CARE formal model through the application of both exhaustive and statistical model checking techniques within Uppaal; (iii) the establishment of a direct connection between the abstract formal model and the actual source code of CARE using traceability and model-based testing. Abstract tests are generated from the Uppaal models and concretised as JUnit tests for the CARE implementation. This connection helps to validate the chosen level of abstraction.
@@ 2. Related work
Other model-based verification approaches for service-oriented systems and distributed middleware are surveyed in [60, 31]. In the literature, there exist many formalisms for modelling and analysing the behaviour of services, ranging from behavioural type systems, including behavioural contracts [29, 1, 49] and session types [26, 42, 34, 28, 55], to automata-based formalisms, including interface automata [3] and (timed) (I/O) automata [54, 5, 33]. Foundational models for service contracts and session types are surveyed in [23, 7, 44]. Sessions and session types [26, 42, 34, 28, 55] have been introduced to reason over the behaviour of services in terms of their interactions. Compared to contract automata, behavioural contracts and session types use process-algebraic frameworks, instead of finite state automata. Indeed, contract automata are similar to other software engineering-based formalisms such as I/O automata [54, 5, 33].
@@ 3.2. Uppaal
Uppaal SMC [32] is a variant of Uppaal [24], which is a well-known toolbox for the verification of real-time systems. Uppaal models are stochastic timed automata, in which non-determinism is replaced with probabilistic choices, and time delays with probability distributions (uniform for bounded time and exponential for unbounded time). These automata may communicate only via broadcast channels and shared variables. In this paper, we will use both exhaustive and statistical model checking. Statistical Model Checking (SMC) [51] involves running a sufficient number of (probabilistic) simulations of a system model to obtain statistical evidence (with a predefined level of statistical confidence) of the quantitative properties to be checked. Monte Carlo estimation with Chernoff-Hoeffding bound executes $N=\lceil(ln(2)-ln(\alpha))/(2\varepsilon^{2})\rceil$ simulations $\rho_{i}$, $i\!\in\!1...N$, to provide the interval $[p^{\prime}-\varepsilon,p^{\prime}+\varepsilon]$ with confidence $1-\alpha$. Here, $p^{\prime}=(\#\{\rho_{i}\!\mid\!\rho_{i}\!\models\!\varphi\})/N$, i.e., $\mathtt{Pr}(|p^{\prime}-p|\leq\varepsilon)\geq 1-\alpha$ where $p$ is the unknown value of the formula $\varphi$ being estimated statistically and $\varepsilon$ and $\alpha$ are the user-defined precision and confidence, respectively. Crucially, the number of simulations used to estimate a formula is independent of the model’s size and depends only on the parameters $\alpha$ and $\varepsilon$. In practice the number of simulations required by Uppaal to reach a specific confidence level is optimized and is thus lower than the above theoretical bound. Uppaal supports template automata used to instantiate different copies (in different experiments) of the same automaton, distinguishable by their parameters. The selection of Uppaal as the chosen tool is influenced by several factors. These include its extensive adoption by the community, expertise of the author, usability, primitive support for real-time and stochastic modelling, probabilistic and non-deterministic choices, and capabilities for statistical model checking, simulation, and model-based testing.
@@
The off-line test case generator of Uppaal is called Yggdrasil, which is integrated since version 4.1 into the Uppaal GUI as a tab. Starting from the Uppaal models, a suite of test cases is created, aiming to cover all syntactic transitions in the model (i.e., edge coverage). The Uppaal model can be decorated with test code, which can be generated from the execution of a transition or at the entry to or exit from a location. Any language can be used for the test code. An abstract test case is generated from a trace of execution of the Uppaal model. During the execution of the trace, the test code of the model is dumped into a text file, forming the abstract test case. When the system is composed of a set of automata, it is possible to generate the test code from a single automaton or from the whole system. In this latter case, the single generated abstract test case contains the interleaved test code from all automata. The traces of execution can be generated as witnesses of some reachability property defined using the Uppaal syntax. Random traces can also be executed to improve test coverage. Parameters to be set include the depth of the trace, the type of search (i.e., breadth, depth), and options for the trace to be generated (i.e., fastest, shortest). The generated abstract test cases need to be decorated and filled with the missing information from the system under test to become concrete test cases. More details will be provided in Section 7.
@@ 3.3. CARE
The Contract Automata Runtime Environment (CARE[15] provides facilities for pairing the contract automata specifications with actual implementations of service-based applications. Note that our purpose is not to formally analyse the applications created using CARE, but to formally analyse CARE itself. The formal verification of applications developed using CARE is a consequence of the reliance of CARE on the formal guarantees provided by contract automata [15], along with the formal verification of CARE as an application. This paper addresses the latter aspect. This software is organised into classes for the orchestrated services (cf. Figure 2) and classes for the orchestrator (cf. Figure 3). The two core abstract Java classes are the RunnableOrchestration and the RunnableOrchestratedContract.
@@
The class diagram of the orchestrator side in Figure 3. The abstract class RunnableOrchestration implements a special service that reads the synthesised orchestration (stored in the instance variable contract) and orchestrates the services (the instances of RunnableOrchestratedContract) to realise the overall application. The instance variables port and addresses contain the corresponding addresses and ports of the services involved in the orchestration.
Choices and Actions
Two aspects to implement for both the orchestrator and the services are choices and termination (through the abstract method choice). Indeed, during the execution of an orchestration, the selection of the transition to execute in the presence of multiple enabled transitions is implemented through this abstract method. CARE is equipped with default implementations, but can be extended (by implementing the relative interfaces and abstract methods) to include other options, other than the default ones. Currently, a so-called ‘dictatorial’ choice (i.e., an internal choice of the orchestrator, external for the services) and a so-called ‘majoritarian’ choice (services vote and the majority wins) are two implemented options. For the services, MajoritarianChoiceRunnableOrchestratedContract and DictatorialChoiceRunnableOrchestratedContract are the two classes specialising RunnableOrchestratedContract according to how the choice is handled and implementing the abstract methods, whilst MajoritarianChoiceRunnableOrchestration and DictatorialChoiceRunnableOrchestration are those specialising RunnableOrchestration. CARE also provides default implementations for the low-level message exchanges. Currently, the two available options are the ‘centralised’ action, where the orchestrator acts as a proxy, and the ‘distributed’ action, where two services matching their actions directly interact with each other once the orchestrator has made them aware of a matching partner and its address/port. Accordingly, each RunnableOrchestratedContract has an OrchestratedAction, and RunnableOrchestration has an OrchestratorAction (instance variable act) used to implement the corresponding actions that can be either distributed or centralised according to the current implementation.
{exa}
We discuss an example of interactions between two services and an orchestrator in CARE. The example comprehends both the CentralisedAction and DistributedAction implementations of CARE. The sequence of service invocations is displayed in Figure 4. In this example, the orchestrator is executing a transition of the orchestration automaton that is labelled with the match [?coffee,!coffee], in which Alice is requesting a coffee and Bob is offering a coffee. As stated above, the RunnableOrchestratedContract class of both Alice and Bob has two instance variables (contract and service, resp.) containing the contract automaton object and the implementation provided as a Java class. In this example, the implementation class of Alice has a method Integer coffee(String arg) that is paired with the corresponding request action ?coffee of her contract. Similarly, the implementation class of Bob has a method String coffee(Integer arg) that is paired with the corresponding offer action !coffee of his contract.
Figure 4. The sequence diagrams depicting the interactions implementing a centralised match action (on the left) and a distributed match action (on the right)
In both centralised and dictatorial configurations, the method coffee of Alice is invoked twice: firstly, passing no argument, it returns an Integer value (e.g., the amount of sugar) that is passed (by the orchestrator ror) as an argument to the method coffee of Bob, which in turn returns a String value that is eventually passed as an argument to the method coffee of Alice, thus fulfilling the coffee request.
In the CentralisedAction implementation, the orchestrator acts as a broker, forwarding to the involved parties the various invocations and responses. In the DistributedAction implementation, the orchestrator communicates the chosen action to both matching services, and the offerer is also notified of a match (rather than an offer) action, so that it can open a fresh port to interact with the requester, which is notified of the address and port of the offerer. Upon successful termination of the interactions, each RunnableOrchestratedContract updates its contract status, and the orchestrator proceeds with the next invocation according to the overall orchestration contract, whose status is also updated. In Section 5 will discuss how this kind of interaction is modelled.
@@ 4. Methodology
In this section, we describe both the methodology used for modelling the communications, abstracting away irrelevant details, and the adequacy of the model to the actual implementation.
@@ 4.1. Modelling TCP/IP sockets communication
The two global (FIFO) arrays in the model are named orc2services[N][queueSize] and services2orc[N][queueSize] (see Figure 5 left). These are used to model the TCP/IP socket buffers, where N is the number of services in the model and queueSize is the size of the queue of the components. These arrays are only modified by functions for enqueueing and dequeuing messages.
@@
The local declarations of the two automata contain methods for sending and receiving from the partners and checking their queues of messages. Both automata declare a method enqueue for sending to the partner a message. Messages are modeled as global integer constants, depicted in Figure 5 left. Indeed, in this model the actual payload of each communication is abstracted away. For example, in Figure 3 left, the constants REQUEST and OFFER abstract payloads that could be any class instantiated by the user, comprehending newly created Java classes (see Section 4.2). The enqueue method of the services only takes as parameter the signal ide_sig to send (see Figure 5 center). The identifier of the service id is only needed on the orchestrator side to identify the partner (there is only one orchestrator thus no identifier is needed for it, see Figure 5 right). Similarly, both automata have a method dequeue for consuming messages from their respective arrays.
@@
Source locations of sending and receiving transitions are neither committed nor urgent. In fact, an enabled committed transition (i.e., whose source location is committed, denoted with C) must be executed before any other non-committed transition in the network. Instead, an urgent transition must be executed without any delay. If a sending or receiving transition were to be either committed or urgent, it could introduce the possibility of false positive deadlocks. This scenario arises when, for example, the receiver has a full buffer (i.e., array) and is prepared to free it, but the sender transition is enabled and committed. Similarly, this false positive can occur when the buffer is full, the send transition is urgent but the receive operation is not urgent, or vice versa, when the buffer is empty, the read transition is urgent, and the send operation is not urgent.
Hence, the operations of writing to and reading from a buffer are represented using stochastic delays, specifically following an exponential distribution. Two rates are employed to capture the delay associated with reading and writing. These exponential delays (variables write and read) are present in all non-committed and non-final locations of both automata in Figure 9 and Figure 10 (modelling the orchestrator and the services, cf. Section 5). However, the presence of unbounded delays introduces scenarios where the receiver (resp., sender) may wait indefinitely without executing its read (resp., write) transition, even when it is enabled. These scenarios would invalidate the exhaustive model checking of reachability properties that are satisfied by the actual system, leading to false positives. In the real implementation, Java TCP/IP sockets offer a timeout mechanism wherein an exception is thrown if no message exchange occurs within a specified time frame. All sockets used by CARE have this timeout. Consequently, a dedicated automaton called SocketTimeout is replicated for each service to model the timeout operation (see Figure 11 left). Each send or receive operation in every socket resets the SocketTimeout clock c. If no reset operation is received within a certain duration (variable timeout), SocketTimeout enters a location called Timeout and broadcasts the signal fail, indicating that an exception has been thrown. All automata have a transition from every location (except for Terminated) to a location Timeout that is reached upon receiving the signal fail. For readability, these Timeout locations and all their incoming fail transitions are not shown in Figure 9 and Figure 10.
@@ 4.2. Abstractions
We have discussed the modelling of Java TCP/IP socket communications. We now discuss other aspects that have been abstracted away in the model. We note that the abstracted aspects are irrelevant for the analysis discussed in Section 6. We remark that the paper objective is to verify the interactions of the CARE middleware itself. Therefore, the underlying application that is executed by CARE is abstracted away. Therefore, in the model, the underlying orchestration automaton is abstracted together with the contracts of the services. Thus, all conditionals that are dependent from the underlying orchestration are abstracted as probabilistic choices. We assume that the orchestration has been correctly synthesised from the services contracts. This allows us to verify the interactions for any possible valid orchestration. If a specific orchestration would be modelled, then we would lose such generality. In particular, the payloads of the communications (e.g., which specific action, which choices) are abstracted. The conditions used to decide whether to perform a choice, an action or to stop are also abstracted away. The only modelled condition is that no two consecutive choices are allowed (i.e., after choosing, the chosen step must be performed). When executing a transition, the conditions used to check whether the label of the transition is an offer or a match are abstracted away in the model. Moreover, the identifiers of the services involved in performing a choice or an action (which are concretely extracted from the labels of the transitions) are also chosen non-deterministically. Indeed, all services are distinguished replicas of the same automaton. Finally, we model a single orchestration. In fact, when multiple orchestrations are executed, they operate independently of each other and can be verified individually.
As stated above, the underlying application executed by CARE is abstracted away. Therefore, properties related to the underlying application are not verifiable under the current abstraction. An example of property not verifiable under the current abstraction could be to show that the returned payload of an action (e.g., an Integer object) is always non-negative, by reasoning on the code of the corresponding method returning the value. This kind of properties cannot be verified under the current abstraction. In this case, a theorem prover like KeY [2] would be more suited than Uppaal.
4.3. Traceability
@@
This paper adopts a lightweight approach to integrate the formal model with the concrete implementation, using the facilities provided by Uppaal. Traceability and model-based testing have been used to ascertain the adequacy of the model with respect to the implementation. Traceability involves connecting the items of the abstract model with the lines of source code that are being abstractly represented. As stated previously, some aspects of the application are abstracted away. Therefore, only the relevant source code lines are connected with the abstract model. In [10] each transition of the model is traced back to the corresponding source code instructions using comments in the model. Traceability information is related to a specific version of the source code [^1] https://github.com/contractautomataproject/CARE/releases/tag/v1.0.1. For example, Figure 6 shows a portion of the automaton modelling the RunnableOrchestratedContract (discussed in details in the next section). Specifically, from state Ready a transition is executed which is traced back to line 98 of the class RunnableOrchestratedContract, also showed in Figure 6. Indeed, at line 98 the first read operation from the ObjectInputStream, reading from the TCP/IP socket, is performed by the service. This is abstracted in the model by the operation dequeue() of the automaton.